# Spin-charge separation and resonant valence bond spin liquid in a frustrated doped Mott insulator

### v1

@@ Abstract
Anderson’s groundbreaking ideas of resonant valence bond (RVB) liquid and spin-charge separation initiated a transformative shift in modern physics. Extensive implications influenced a broad spectrum of fields, from high-temperature superconductors to quantum computing, and gave birth to key concepts in physics, such as quantum spin liquids, emergent gauge symmetries, topological order, and fractionalisation. Despite extensive efforts to demonstrate the existence of an RVB phase in the Hubbard model, a definitive realisation has proven elusive. Here, we present a concise, realistic, and elegant solution to this longstanding problem by demonstrating analytically that an RVB spin liquid, exhibiting spin-charge separation, emerges as the ground state of doped Mott insulators on corner-sharing tetrahedral lattices with frustrated hopping near half-filling – a manifestation of the counter-Nagaoka effect. We confirm numerically that this result holds for finite-size systems, finite dopant density, and small exchange interactions. While much attention has been devoted to the emergence of new states from geometrically frustrated interactions, our work demonstrates that kinetic energy frustration in doped Mott insulators may be pivotal to stabilise robust, topologically ordered states in real materials.
Inspired by Pauling’s early theory of resonant covalent-bond-sharing in aromatic molecules, such as benzene, and by Bethe’s antiferromagnetic linear chain, Fazekas and Anderson [Anderson1973, Fazekas1974] proposed in the early 1970’s a possible competitor to Néel order in antiferromagnetically coupled spins: a state where spins are valence bond paired into singlets that resonate quantum mechanically amongst different pairing configurations, thus compensating for the partial loss of antiferromagnetic energy – forming a resonant valence bond (RVB) state. The exciting idea gained further momentum with the discovery of high-temperature superconductivity about a decade later, and with Baskaran and Anderson’s prompt proposal [Anderson1987, Anderson1987b, Anderson1997] that it may indeed be underpinned by spin-charge separation [Baskaran1987, Zhu1988, Anderson1987b, Kivelson1987] in a system of strongly correlated electrons in an RVB liquid phase (for a review, see Refs. Anderson2004, Lee2006). The appeal was so great that for decades to follow numerous physicists have been looking for evidence of this suggestion “along a bewildering variety of routes” [Anderson2004].
@@
Figure 1: Pictorial illustration of one of the dimer-singlet (dark thin ellipsoids) states with a holon at a given location (light sphere), that participate in the superposition forming the RVB liquid ground state of the $t$-$J$ model, Eq. (1), on a lattice of corner-sharing tetrahedra. The $\pi$-fluxes depict the phase induced by the singlet coverings (relative to a ferromagnetic background) when the holon moves around the corresponding triangular plaquette. The inset shows an Y2Ir2O7 pyrochlore crystal grown by D. Prabhakaran in Oxford.
@@
Intense work continued throughout the years, yet Anderson’s original proposal has hitherto evaded discovery. Following the seminal work by Moessner and Sondhi [Moessner2001], which established the possibility of a short-ranged liquid phase in quantum dimer models, several efforts attempted to translate this result into a system of SU(2) spins, building on Klein models [Klein1982], and effective and decorated dimer models [Batista2004, Raman2005, Normand2014, Normand2016]. It is generally believed that realistic Hamiltonians harbouring RVB spin liquid phases and spin-charge separation must exist, but to date they have remained elusive.
@@
In this work, we find a concise and elegant solution – unintentionally fulfilling the inspirational words of Ref. Raman2005: “a more important task perhaps, now that the question of principle is settled, is to refocus on studying much simpler Hamiltonians”. We demonstrate that an RVB liquid phase, exhibiting spin-charge separation, emerges naturally upon doping a large-$U$ Hubbard model on a lattice of corner-sharing tetrahedra – be it the planar checkerboard or the three-dimensional pyrochlore lattice – near half-filling. This result hinges on kinematic spin correlations set by frustrated hopping and is an instance of the so-called ‘counter-Nagaoka effect’ [Haerter2005]. Without loss of generality, in the following we shall consider the case of hole doping; the same can be said for electron doping when the hopping amplitude has the opposite sign. We prove our claims by finding an exact lower bound to the Hamiltonian energy, and by explicitly constructing an RVB liquid state that precisely matches it in the thermodynamic single-hole limit. We confirm this behaviour by investigating numerically systems of finite size. We propose a way to construct single-hole and two-holes RVB liquid ground state wave functions for any size that appear to be exact within numerical accuracy – a truly remarkable feat for a strongly correlated electron problem. Finally, we show numerically that our results are stable in systems of finite hole density, as well as in presence of sufficiently small exchange interactions between the spins [footnote:Normand].
@@
The crucial role of kinetic energy frustration in stabilising a topologically ordered RVB liquid state is a remarkable outcome of our findings, which provide a novel way to connect dimer coverings with solutions to the Hubbard model. Our results have the potential to shift the traditional paradigm, which has predominantly considered exchange frustration as the mechanism for stabilising quantum spin liquids. In other words, doping insulators that lie deep within the Mott regime with a single type of carrier (electrons or holes, depending on the sign of the hopping amplitude) emerges as a new guiding principle for the discovery of spin liquid states in correlated matter.
@@
Our Hamiltonian is eminently simple and realistic, and of direct relevance to several experimental settings as well as synthetic quantum platforms [Feynman1982, Trabesinger2012, Georgescu2014]. We identify in particular some families of pyrochlore compounds as suitable frameworks to experimentally test our predictions and suggest how it may be possible to manifest the elusive RVB state for the first time in these materials.
@@ I Results
We study the $t$-$J$ model on a lattice of corner-sharing tetrahedra [footnote:checkerboard] (see Fig. 1) near half-filling, with Hamiltonian
^$\hat{\mathcal{H}}=-t\sum_{\langle i,j\rangle\;\sigma}\left[\hat{c}^{\dagger}_{i\sigma}\hat{c}_{j\sigma}+h.c.\right]+J\sum_{\langle i,j\rangle}\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j}\,,$ (1)
where $\langle i,j\rangle$ denotes nearest-neighbour pairs of sites; $\hat{c}^{\dagger}_{i}$ ($\hat{c}_{i}$) is the constrained fermionic creation (annihilation) operator at site $i$, acting in the space where double occupancy is strictly forbidden (see Supp. Mat. Note C); and $\hat{\bm{S}}_{i}$ is the corresponding spin operator, $\hat{\bm{S}}_{i}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{i\sigma}{\bm{\sigma}}_{\sigma\sigma^{\prime}}\hat{c}_{i\sigma^{\prime}}$. Here $\sigma$ labels the states of spin-$1/2$ degrees of freedom, and ${\bm{\sigma}}$ is the conventional vector of Pauli matrices. For generality, we treat the hopping $t$ and the spin exchange $J$ as independent parameters. Note that for fermions and our convention in Eq. (1), hole doping with $t>0$ or electron doping with $t<0$ correspond to frustrated hopping when the elementary loop of the lattice has an odd number of sites. In the following we shall focus, without loss of generality, on the case of hole doping and $t>0$.
Our first result is analytical, derived in the thermodynamic limit for a single hole and $J=0$. As demonstrated in detail in the Supp. Mat. Note A, the spectrum of the Hamiltonian in Eq. (1) has a strict lower bound of $-4t$. Furthermore, we prove that this bound is met by the RVB spin liquid ground state
@@
where ${\bm{r}}$ spans all possible positions of the holon; $a_{\bm{r}}$ is a generic dimer covering of the lattice with a holon at ${\bm{r}}$ subject to the condition that there is one and only one dimer per tetrahedron (see e.g., Fig. 1); and $\hat{d}^{\dagger}_{ij}=\frac{1}{\sqrt{2}}(\hat{c}^{\dagger}_{i\uparrow}\hat{c}^{\dagger}_{j\downarrow}-\hat{c}^{\dagger}_{i\downarrow}\hat{c}^{\dagger}_{j\uparrow})$. We shall refer to each state in the superposition forming $|\Psi_{\rm RVB}\rangle$ as a dimer-singlet state. The kinetic energy frustration disappears for the opposite sign of the hopping amplitude, which is equivalent to inserting a $\pi$-flux on each triangular plaquette of the pyrochlore lattice. The RVB wave function minimises the holon kinetic energy by spontaneously generating a $\pi$-flux on the triangular face containing the hole and a singlet. Namely, the RVB wave function binds two $\pi$-fluxes (fluctuating with the dimer-singlet resonant states) to each hole to lower its kinetic energy to $-4t$ by partially lifting the frustration (see Fig. 1).
This is an instance of the so-called counter-Nagaoka effect [Haerter2005], whereby frustrated quantum hole hopping leads to antiferromagnetic correlations in a Mott insulator near half-filling. While the phenomenon has been shown to generically lead to conventionally ordered phases in two dimensions [Sposetti2014, Lisandrini2017, Zhang2018, Morera2021, Kim2023], it had not been thus far investigated in three dimensions, and it is all the more remarkable to see it favour a quantum spin liquid state in lattices of corner-sharing tetrahedra (for a brief review on the counter-Nagaoka effect, see Supp. Mat. Note B).
Figure 2: Top panel: Spin structure factor for the 32-site system with one hole (top half) and two holes (bottom half), for $t=1$ and $J=0$. Accessible momenta are marked by grey pluses. Bottom panel: Blue dots show the holon density for a 16-site and a 32-site system doped with a single hole as a function of distance from the pinned spin [footnote1], indicating that the spinon and holon separate. Crosses show the holon-holon correlations when there are two holes present in the system, and they are numerically identical to the holon-spinon correlations.
@@
We studied the behaviour of Eq. (1) in finite systems using exact diagonalisation (ED) on a 16-site cubic cell, and density matrix renormalisation group (DMRG) [White1992, Schollwock2005, Schollwock2011] on a 32-site system, with periodic boundary conditions (for details of the study, see Supp. Mat. Note D). Remarkably, we find that the ground state energy bound $-4t$ is met (when $J=0$) for all system sizes, within numerical precision. A non-trivial degeneracy is observed ($18$ states for 16 sites and $34$ states for 32 sites). Consistently with the analytical considerations above, we find a unique ground state (again with energy $-4t$) when the unpaired spin is pinned to a given lattice site.
@@
Within numerical accuracy, we show that the ground state wave function with a fixed spin is given by the equal amplitude superposition of all dimer-singlet configurations (constrained to no more than one dimer per tetrahedron) with a delocalised holon, for our 16-site and 32-site systems (see Supp. Mat. Note D). This allows us to propose an exact wave function for the single-hole case for all system sizes – which is a highly unusual feat. Such a state has vanishing expectation value of the projector onto the maximum angular momentum for each tetrahedron with no hole [Kivelson_unpub, Batista2004, Raman2005], which we verified to be the case (Supp. Mat. Note D).
@@
As illustrated in Fig. 2, the holon density shows weak repulsive dependence on the distance from the spinon in finite size systems, in agreement with spin-charge separation in the thermodynamic limit, argued earlier on analytical grounds (see also Supp. Mat. Note D). In the top panel of Fig. 2, we show the spin structure factor of our numerical ground states. Characteristic differences that could be measured experimentally seemingly distinguish it from, for example, the classical or quantum Heisenberg spin liquid structure factors (see Supp. Mat. Note D); however future work on larger system sizes is needed to make more precise statements about this. On general grounds, we expect exponentially decaying spin correlations; the nature of the singlet correlations is less clear, as the correspondence with spin ice states [Baskaran1988, Read1989, Fradkin1990, Raman2005] could induce power law behaviour and an emergent $U(1)$ gauge symmetry with gapless photon excitations.
[…]
@@
Intriguingly, the opposite (stoichiometric) limit of vanishing hole density is singular for the 16-site and 32-site systems in question. The wave function given by the equal amplitude superposition of all dimer-singlet states with one and only one dimer per tetrahedron vanishes identically [footnote:vanishing] (see Supp. Mat. Note A). The RVB liquid phase appears to exist solely in presence of hole doping (at least within the remit of our accessible system sizes and periodic boundary conditions), and therefore it can only be stabilised by kinetic energy frustration.
@@
Finally, we turn to look at the concomitant effects of $t$ and $J$ in the system. We consider the case of a 16-site system with a single hole; we fix $t=1$ and vary $J$, spanning both positive and negative values. In Fig. 3, we plot the behaviour of the ground state spin structure factor for all distinct points in reciprocal space. We find that the spin liquid behaviour persists down to a finite negative value of $(J/t)_{c}\simeq-0.052$ (for holon density $6.25$%), suggesting that a transition from ferromagnet to quantum spin liquid (QSL) may be induced on the pyrochlore lattice by frustrated hole doping. The effective doping-induced antiferromagnetic interaction between the spins has been argued to be proportional to the holon density at sufficiently small concentrations [Haerter2005], and we expect our results to correspondingly become more pronounced [footnote:caveat]. Indeed, adding a second hole to the 16-site system, we find $(J/t)_{c}\simeq-0.173$ (holon density $12.5$%), see Supp. Mat. Note E. In Fig. 3, we further compare for reference the case of electron doping, which is not frustrated for $t>0$ and exhibits a kinetic tendency towards ferromagnetic order through the Nagaoka effect.
@@
Figure 3: Left panel: Dependence of the ground state spin structure factor values (for all individual reciprocal lattice points) on $J/t$ ($t>0$) for a 16-site system with one hole. The region where the correlations are antiferromagnetic in spite of ferromagnetic $J$ is shaded. Right panel: same data when the system is doped with one electron, for comparison.

### v2

# A resonant valence bond spin liquid in the dilute limit of doped frustrated Mott insulators

@@ Abstract
Ideas about resonant valence bond liquids and spin-charge separation have led to key concepts in physics such as quantum spin liquids, emergent gauge symmetries, topological order, and fractionalisation. Despite extensive efforts to demonstrate the existence of a resonant valence bond phase in the Hubbard model that originally motivated the concept, a definitive realisation has yet to be achieved. Here we present a solution to this long-standing problem by uncovering a resonant valence bond phase exhibiting spin-charge separation in realistic Hamiltonians. We show analytically that this ground state emerges in the dilute-doping limit of a half-filled Mott insulator on corner-sharing tetrahedral lattices with frustrated hopping, in the absence of exchange interactions. We confirm numerically that the results extend to finite exchange interactions, finite-sized systems and finite dopant density. Although much attention has been devoted to the emergence of unconventional states from geometrically frustrated interactions, our work demonstrates that kinetic energy frustration in doped Mott insulators may be essential for stabilising robust, topologically ordered states in real materials.
Inspired by Pauling’s early theory of resonant covalent-bond-sharing in aromatic molecules, such as benzene, and by Bethe’s antiferromagnetic linear chain, Fazekas and Anderson [Anderson1973, Fazekas1974] proposed in the early 1970s a possible competitor to Néel order in antiferromagnetically coupled spins: a state where spins are valence bond paired into singlets that resonate quantum mechanically amongst different pairing configurations, thus compensating for the partial loss of antiferromagnetic energy – forming a resonant valence bond (RVB) state. The exciting idea gained further momentum with the discovery of high-temperature superconductivity about a decade later, and with the prompt proposal that it may indeed be underpinned by spin-charge separation [Kivelson1987, Anderson1987, Baskaran1987, Anderson1987b, Zhu1988, Anderson1997] in a system of strongly correlated electrons in an RVB liquid phase (for a review, see Refs. Anderson2004, Lee2006). The appeal was so great that for decades to follow numerous physicists have been looking for evidence of this suggestion “along a bewildering variety of routes” [Anderson2004].
@@
Figure 1: Pictorial illustration of one of the dimer-singlet states. These states participate in the superposition forming the RVB liquid ground state of the $t$-$J$ model, Eq. (1), on a lattice of corner-sharing tetrahedra. Singlets are indicated by dark thin ellipsoids, and a holon at a given location by a light sphere. The $\pi$-fluxes depict the phase induced by the singlet coverings (relative to a ferromagnetic background) when the holon moves around the corresponding triangular plaquette. The inset shows an Y2Ir2O7 pyrochlore crystal grown by D. Prabhakaran in Oxford.
@@
Intense work continued throughout the years, yet Anderson’s original proposal has hitherto evaded discovery. Following the seminal work by Moessner and Sondhi [Moessner2001], which established the possibility of a short-ranged liquid phase in quantum dimer models [Rokhsar1988], several efforts attempted to translate this result into a system of SU(2) spins, building on Klein models [Klein1982], and effective and decorated dimer models [Batista2004, Raman2005, Normand2014, Normand2016]. It is believed that magnetic frustration in quasi-two-dimensional (quasi-2D) Mott insulators could be responsible for RVB physics; however, parent Hamiltonians that host an exact RVB ground state are often unphysical (see e.g., Ref. Schuch2012).
@@
In this work, we find a concise and elegant solution – unintentionally fulfilling the inspirational words of Ref. Raman2005: “a more important task perhaps, now that the question of principle is settled, is to refocus on studying much simpler Hamiltonians”. We demonstrate that an RVB liquid phase, exhibiting spin-charge separation, emerges naturally upon dilute-doping a large-$U$ (on-site Coulomb repulsion) Hubbard model on a lattice of corner-sharing tetrahedra – be it the planar checkerboard or the three-dimensional pyrochlore lattice – near half-filling. This result hinges on kinematic spin correlations set by frustrated hopping and is an instance of the so-called ‘counter-Nagaoka effect’ [Haerter2005]. Without loss of generality, in the following we shall consider the case of hole doping; the same can be said for electron doping when the hopping amplitude has the opposite sign. We prove our claims by finding an exact lower bound to the Hamiltonian energy, and by explicitly constructing an RVB liquid state that precisely matches it in the thermodynamic single-hole limit. We confirm this behaviour by investigating numerically systems of finite size. (A numerical study of one- and two-hole doping in the $t$-$J$ model on the checkerboard lattice had already been carried out by Poilblanc [Poilblanc2004], where $t$ is the electron hopping amplitude and $J$ the nearest-neighbour exchange interaction strength. Poilblanc’s work highlighted the 4t-4t and 8t-8t respective ground state energies, and arguing how the singlet background somehow changes the effective hopping – a precursor of the understanding in terms of RVB liquid and two $\pi$-flux attachment presented in this work.) Here we propose a way to construct single-hole and two-holes RVB liquid ground state wave functions for any system size. These seem to be exact within numerical accuracy, a truly notable feat for a strongly correlated electron problem. Finally, we show numerically that our results are stable in systems of finite hole density, as well as in presence of sufficiently small exchange interactions between the spins.
@@
The crucial role of kinetic energy frustration in stabilising a topologically ordered RVB liquid state is a remarkable outcome of our findings, which provide a hitherto-unexplored way to connect dimer coverings with solutions to the Hubbard model. Our results have the potential to shift the traditional paradigm, which has predominantly considered exchange frustration as the mechanism for stabilising quantum spin liquids. In other words, doping insulators that lie deep within the Mott regime with a single type of carrier (electrons or holes, depending on the sign of the hopping amplitude) promises to become a guiding principle for the discovery of spin liquid states in correlated matter.
@@
Our Hamiltonian is eminently simple and realistic, and of direct relevance to several experimental settings as well as synthetic quantum platforms [Feynman1982, Trabesinger2012, Georgescu2014]. We identify in particular some families of pyrochlore compounds as suitable frameworks to experimentally test our predictions and suggest how it may be possible to manifest the elusive RVB state in these materials.
@@ I Results
We study the $t$-$J$ model on a lattice of corner-sharing tetrahedra (see Fig. 1) near half-filling, with Hamiltonian
$\hat{\mathcal{H}}=-t\sum_{\langle i,j\rangle\;\sigma}\left[\hat{c}^{\dagger}_{i\sigma}\hat{c}_{j\sigma}+h.c.\right]+J\sum_{\langle i,j\rangle}\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j}\,$, (1)
where $\langle i,j\rangle$ denotes nearest-neighbour pairs of sites; $\hat{c}^{\dagger}_{i}$ ($\hat{c}_{i}$) is the constrained fermionic creation (annihilation) operator at site $i$, acting in the space where double occupancy is strictly forbidden (see Supp. Mat. Note C); and $\hat{\bm{S}}_{i}$ is the corresponding spin operator, $\hat{\bm{S}}_{i}=\tfrac{1}{2}\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{i\sigma}{\bm{\sigma}}_{\sigma\sigma^{\prime}}\hat{c}_{i\sigma^{\prime}}$. Here $\sigma$ labels the states of spin-$1/2$ degrees of freedom, and $\bm{\sigma}$} is the conventional vector of Pauli matrices. For generality, we treat the hopping $t$ and the spin exchange $J$ as independent parameters. Note that for fermions and our convention in Eq. (1), hole doping with $t>0$ or electron doping with $t<0$ correspond to frustrated hopping when the elementary loop of the lattice has an odd number of sites. In the following we shall focus, without loss of generality, on the case of hole doping and $t>0$.
Our results apply straightforwardly, e.g., to the 2D case of the checkerboard lattice, or to the three-dimensional case of the pyrochlore lattice. Note that for the 16-site system considered below, the two lattices in fact coincide.
Our first result is analytical, derived in the thermodynamic limit for a single hole and $J=0$. As demonstrated in detail in the Supp. Mat. Note A, the spectrum of the Hamiltonian in Eq. (1) has a strict lower bound of 4t-4t. Furthermore, we prove that this bound is met by the RVB spin liquid ground state
@@
where $\bm{r}$} spans all possible positions of the holon; $a_{\bm{r}}$ is a generic dimer covering of the lattice with a holon at $\bm{r}$} subject to the condition that there is one and only one dimer per tetrahedron (see e.g., Fig. 1); and $\hat{d}^{\dagger}_{ij}=\frac{1}{\sqrt{2}}(\hat{c}^{\dagger}_{i\uparrow}\hat{c}^{\dagger}_{j\downarrow}-\hat{c}^{\dagger}_{i\downarrow}\hat{c}^{\dagger}_{j\uparrow})$. We shall refer to each state in the superposition forming $|\Psi_{\rm RVB}\rangle$ as a dimer-singlet state. The kinetic energy frustration disappears for the opposite sign of the hopping amplitude, which is equivalent to inserting a $\pi$-flux on each triangular plaquette of the pyrochlore lattice. The RVB wave function minimises the holon kinetic energy by spontaneously generating a $\pi$-flux on the triangular face containing the hole and a singlet. Namely, the RVB wave function binds two $\pi$-fluxes (fluctuating with the dimer-singlet resonant states) to each hole to lower its kinetic energy to 4t-4t by partially lifting the frustration (see Fig. 1[Poilblanc2004].
This is an instance of the so-called counter-Nagaoka effect [Haerter2005], whereby frustrated quantum hole hopping leads to antiferromagnetic correlations in a Mott insulator near half-filling. The phenomenon has been shown to generically lead to conventionally ordered phases in two dimensions [Sposetti2014, Lisandrini2017, Zhang2018, Morera2024], with the notable exception of a Husimi cactus of triangles (akin locally to a kagome lattice) where Kim proved that it leads to a regular valence bond pattern in presence of a single (deconfined) hole [Kim2023]. It had not been thus far investigated in three dimensions, and it is all the more remarkable to see it favour a quantum spin liquid state in lattices of corner-sharing tetrahedra (for a brief review on the counter-Nagaoka effect, see Supp. Mat. Note B).
Note that, unlike the triangular case considered in Ref. Kim2023, where the dimer-singlet configuration is fully determined once the hole position is fixed, for our corner-sharing tetrahedra there are extensively many dimer-singlets forming a massively entangled superposition of electronic states for each fixed hole position. The latter is exactly what one expects of an RVB liquid state.
Figure 2: Correlations and spin structure factor in the RVB state. Top, blue dots show the holon density $\langle\hat{O}\rangle=\langle 1-\hat{n}_{i}\rangle$ ($\hat{n}_{i}$ being the electron density operator) for a 16-site and a 32-site pyrochlore system doped with a single hole as a function of distance from the pinned spin, $|\bm{r}_{i}-\bm{r}_{0}|$, for system sizes N=16, 32N=16,\,32. It indicates that the spinon and holon separate (the two different values at distance $\sim 0.7$ are due to the two inequivalent third-nearest-neighbour sites on the pyrochlore lattice). Crosses show the holon-holon correlations when there are two holes present in the system. They are numerically identical to the holon-spinon correlations. Bottom, spin structure factor in the $(hhk)$ and $(hk0)$ reciprocal lattice planes for the 32-site pyrochlore system with one hole (top half) and two holes (bottom half), for $t=1$ and $J=0$. Accessible momenta are marked by grey pluses.
@@
We studied the behaviour of Eq. (1) in finite pyrochlore systems using exact diagonalisation (ED) on a 16-site cubic cell, and density matrix renormalisation group (DMRG) [White1992, Schollwock2005, Schollwock2011] on a 32-site system, with periodic boundary conditions (for details of the study, see Supp. Mat. Note D). Remarkably, we find that the ground state energy bound 4t-4t is met (when $J=0$) for all system sizes, within numerical precision. A non-trivial degeneracy is observed ($18$ states for 16 sites and $34$ states for 32 sites). Consistently with the analytical considerations above, we find a unique ground state (again with energy 4t-4t) when the unpaired spin is pinned to a given lattice site.
@@
Within numerical accuracy, we show that the ground state wave function with a fixed spin is given by the equal amplitude superposition of all dimer-singlet configurations (constrained to no more than one dimer per tetrahedron) with a delocalised holon, for our 16-site and 32-site systems (see Supp. Mat. Note D). This allows us to propose an exact wave function for the single-hole case for all system sizes – which is a highly unusual feat. Such a state has vanishing expectation value of the projector onto the maximum angular momentum for each tetrahedron with no hole [Batista2004, Raman2005] (see also Kivelson, S.A., private communication, unpublished), which we verified to be the case (Supp. Mat. Note D).
@@
As illustrated in the top panel of Fig. 2, the holon density shows weak repulsive dependence on the distance from the spinon in finite size systems, in agreement with spin-charge separation in the thermodynamic limit, argued earlier on analytical grounds (see also Supp. Mat. Note D). In the bottom panel of Fig. 2, we show the spin structure factor of our numerical ground states. Characteristic differences that could be measured experimentally seemingly distinguish it from, for example, the classical or quantum Heisenberg spin liquid structure factors (see Supp. Mat. Note D); however future work on larger system sizes is needed to make more precise statements about this. On general grounds, we expect exponentially decaying spin correlations; the nature of the singlet correlations is less clear, as the correspondence with spin ice states [Baskaran1988, Read1989, Fradkin1990, Raman2005] could induce power law behaviour and an emergent U(1) gauge symmetry with gapless photon excitations; this is however far from trivial due to the non-orthogonality of the dimer-singlet states.
We briefly studied the case $J=0$ and two holes, which we were able to access numerically in the 16-site and 32-site systems. We observe no tendency towards holon attraction (see top panel of Fig. 2 and Supp. Mat. Note E). The ground state is now unique, with energy 8t-8t, suggesting that the second holon is able to occupy the site of the unpaired spin, and to render it itinerant irrespective of the presence of the other holon. Consistently, we find that the holon-spinon correlations for a single hole are numerically identical to the holon-holon correlations for two holes (Fig. 2). The nature of the two-holes ground state can be understood from the perspective of the $\pi$-flux binding discussed earlier. A holon can only move on a tetrahedron around the triangular face that contains a singlet (see Fig. 1), and it is forbidden by interference effects to visit the fourth tetrahedral site (thus rendering two holons on the same tetrahedron blind to one another). In other words, the single-hole RVB wave function is an equal weight superposition of the holon propagating in all possible sublattices that can be constructed by choosing one triangle of each tetrahedron that is visited. For the two-holes RVB wave function (namely, the equal amplitude superposition of all dimer-singlet states hosting two holons, with no more than one dimer per tetrahedron), the holons propagate through different triangles when they meet at the same tetrahedron, meaning that the effective coordination number is still equal to 4 for each of them, and the energy is $2\times(-4t)=-8t$. These observations, in particular the coincidence of the spinon-holon and holon-holon correlators in Fig. 2, are strongly suggestive that the holon excitations in our RVB state are bosonic.
We verified numerically that the proposed RVB wave function is also the exact ground state for the two-holes system (for 16 sites and 32 sites). The argument does not extend to three or more holes on finite systems, and their behaviour is left to future investigations.
@@
Intriguingly, the opposite (stoichiometric) limit of vanishing hole density is singular for the 16-site and 32-site systems in question. The wave function given by the equal amplitude superposition of all dimer-singlet states with one and only one dimer per tetrahedron vanishes identically (see Supp. Mat. Note A). We note that the vanishing occurs also for equal amplitude superpositions of all dimer-singlet states, without the constraint of one dimer per tetrahedron — a curious fact that, to the best of our knowledge, had not been pointed out in the literature to date. The RVB liquid phase therefore appears to exist solely in presence of hole doping (at least within the remit of our accessible system sizes and periodic boundary conditions), and it can only be stabilised by kinetic energy frustration.
@@
Finally, we turn to look at the concomitant effects of $t$ and $J$ in the system. We consider the case of a 16-site system with a single hole; we fix $t=1$ and vary $J$, spanning both positive and negative values. In Fig. 3, we plot the behaviour of the ground state spin structure factor for all distinct points in reciprocal space. We find that the spin liquid behaviour persists down to a finite negative value of $(J/t)_{c}\simeq-0.052$ (for holon density 6.256.25%), suggesting that a transition from ferromagnet to quantum spin liquid (QSL) may be induced on the pyrochlore lattice by frustrated hole doping. The effective doping-induced antiferromagnetic interaction between the spins has been argued to be proportional to the holon density at sufficiently small concentrations [Haerter2005], and we expect our results to correspondingly become more pronounced (with the caveat of possible phase separation effects as in Ref. Eisenberg2002 for the large but finite $U$ Hubbard limit). Indeed, adding a second hole to the 16-site system, we find $(J/t)_{c}\simeq-0.173$ (holon density 12.512.5%), see Supp. Mat. Note E. In Fig. 3, we further compare for reference the case of electron doping, which is not frustrated for $t>0$ and exhibits a kinetic tendency towards ferromagnetic order through the Nagaoka effect.
@@
Figure 3: Spin structure factor. Dependence of the ground state spin structure factor $S(\bm{Q})$ (for all individual reciprocal lattice points $\bm{Q}$) on $J/t$ ($t>0$) for a 16-site system. Left, the system is doped with one hole. The region where the correlations are antiferromagnetic, despite the ferromagnetic $J$, is shaded. Right, the system is doped with one electron, for comparison.